OpenStudy (anonymous):

how do you factor fractions? (1/(x+h))(1/(x+h))=?

OpenStudy (anonymous):

\[(\frac{ 1 }{ (x+h) })^{2}\]or\[(\frac{ 1 }{ (x+h)})(\frac{ 1 }{ (x+h) })\]

OpenStudy (anonymous):

either way you write it, how do i solve this

OpenStudy (campbell_st):

to solve it, it needs to be equal to something. All you have so far is an expression

OpenStudy (anonymous):

isnt it something like cross multiplying or something along those lines?

OpenStudy (anonymous):

i just need to factor it out, this is a baby part of my over all problem

OpenStudy (campbell_st):

1st identify the restriction, what value can't x take. if its equal to a value then multiply both sides of the equation by the denominator \[\frac{1}{(x + h)^2} = 4...then..... 1 = 4 \times (x + h)^2\] it looks a lot like 1st principles...

OpenStudy (anonymous):

ahh got it! thanks